Q4. Find the support reaction at A and B 500 N 4 m A 30°
Answer #1

500N -4m A RAR 300 500N 3m Ray 1.15% B Box Roa 01 104 m Go Rey → Rot Rox = Rocos 60 RECOS 60 = 0.5 RB Rey - Rgsingi - 0.866 R- E Fy=0 Ray - Ray-(4/3) 500 500 =0 RAY-0.866R3 = 400 2 EMA. - )(500) C4) – Rey (4-3-6) +Rgi (1-5) =0 - 1600- 0.866 RG (104)Fromm equations 0) Rax - 300-0.5RB Rox=300-0-55(-3460.2) Ron- = 300+ 1730.1 Rog- 2030.1 N